OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. x + y = 3 2x + 2y = 10

mathslover (mathslover):

\[\huge{x+y=3}\] \[\huge{x=3-y}\] 2x+2y = \(\huge{2(3-y)+2y=6-2y+2y = 6 }\) hence no solution

mathslover (mathslover):

note that 2x+2y = 2(x+y) = 2(3) given x + y =3 hence 6 = 2x+2y but given 2x+2y = 10 that is not possible

OpenStudy (anonymous):

not possible...ansr cms only with slope form